# Your copula is a classifier in disguise: classification-based copula density estimation

### v1

@@ 1 Introduction
Figure 1: Our model estimates copula densities through classification. First panel: dependent copula samples. Second panel: we transform the dependent copula samples to a Gaussian scale and train a classifier to distinguish between dependent samples and an independent Gaussian noise. Third panel: this classifier learns to identify dependence and ultimately recovers the copula density, shown on the log scale above.
Copulas are a popular tool for modelling multivariate densities by decomposing the estimation procedure into two steps: first, learning the marginal density of each variable; and second, stitching those marginals together through a copula model. This adoption of copulas is primarily due to their distinct advantages over other models as a consequence of this two-step procedure. Namely, marginal models can be purpose-built for each dimension, which is relevant for applications, e.g. in neuroscience (Berkes et al.,, 2008), climate research (Huk et al.,, 2023), econometrics (Czado et al.,, 2022), and domain adaptation (Lopez-Paz et al.,, 2012). Further, copulas focus all optimisation efforts uniquely on modelling dependence, leading to higher quality densities and samples (Tagasovska et al.,, 2019), and allows for the estimation of marginal models in parallel or in real time as used in Bayesian density estimation (Huk et al.,, 2024) and reinforcement learning (Wang et al.,, 2021).
@@
To this aim, we leverage a novel connection between copula densities and classification to propose a new copula estimation framework. This involves training a classifier to distinguish observed dependent data from an independent version of itself, recovering the true copula density at optimality. After providing a background on copula density estimation and probabilistic classification for density ratios in Section 2, we elicit the connection between copulas and classification in Section 3 and develop our classification-based copula estimation scheme. To the best of our knowledge, this is the first methodology bridging copula estimation with probabilistic classification. We showcase equivalences between well-known copulas and their classifier counterparts in Section 4. We empirically validate our approach in Section 5 by training copulas using classification and demonstrating their superior density estimation and sampling capabilities compared to relevant benchmarks. Finally, we discuss related works in Section 6 and conclude in Section 7.
@@ 2 Background
Let $p(\mathbf{x})$ be a multivariate density over $\mathcal{X}\subseteq\mathbb{R}^{d}$, from which we observe i.i.d. samples $\mathcal{D}_{P}=\{\mathbf{x}^{k}\}_{k=1}^{T_{p}}\sim p(\mathbf{x})$. Similarly, let $p_{1}(x_{1}),\ldots,p_{d}(x_{d})$ be the marginal densities of $p(\mathbf{x})$, over $\mathcal{X}_{i}\subseteq\mathbb{R}$ with corresponding i.i.d. samples $\mathcal{D}_{P_{i}}=\{{x_{i}}^{k}\}_{k=1}^{T_{q}}\sim p_{i}({x_{i}})$ for $i=1,\ldots,d$, and assume that all distributions are continuous. Further, let $P$ and $P_{1},\ldots,P_{d}$ be the respective cumulative distribution functions (CDFs) of the previously mentioned densities, in order of appearance. We define applications of CDFs such as $\mathcal{P}(\mathbf{x})\equiv(P_{1}(x_{1}),\ldots,P_{d}(x_{d}))$ as shorthand for the dimension-wise application of marginal CDFs to a vector or $\mathcal{P}(\mathcal{D})$ to mean their application to every vector element of a set. Probability and cardinality of a set $A$ are denoted by $\mathbb{P}(A)$ and $|A|$ correspondingly.
@@ 2.1 Copula-Based Density Estimation
Sklar’s theorem (Sklar, (1959) found in Appendix B) states that any joint probability density function with continuous marginal densities can always be uniquely factorised into the product of its marginals and a copula density $c:[0,1]^{d}\mapsto(0,\infty)$ in the following way:
@@
Consequently, one can decompose the estimation problem of learning $p(\mathbf{x})$ into first learning all marginals $\{p_{i}\}_{i=1}^{d}$, and then inferring an appropriate copula model for the dependence structure.
@@
There exist many types of copulas, both parametric and non-parametric. Most parametric copulas are only really suited for 2D dependence models and greatly suffer from the curse of dimensionality (Hofert et al.,, 2013). The Gaussian copula is a popular parametric choice as it can be fitted quickly even to moderate dimensions. However, it lacks the desired flexibility to capture complex patterns. Among non-parametric copulas, kernel density estimator (KDE) copulas (Geenens et al.,, 2017) are commonly used. They transform the observed data via appropriate (often Gaussian) inverse CDFs to a latent space and perform regular KDE on the latent density. However, this KDE copula method suffers from drawbacks common to KDE methods in higher dimensions. Deep learning copula models are still a nascent line of research, with a handful of candidate solutions such as Kamthe et al., (2021); Janke et al., (2021); Ashok et al., (2024). The Implicit Generative Copula (IGC) of Janke et al., (2021) is a promising model, which works by generating data in a latent space and applying an approximate integral transformation to recover samples in $[0,1]$. As an implicit model only able to generate samples, the IGC cannot evaluate densities, thus preventing many applications of copula models. As such, current copula models supporting density estimation and sampling are mostly limited to low or medium-dimensional modelling.
@@
The state-of-the-art approach is vine copula models, offering a divide-and-conquer solution to expand copula models to higher dimensions. Vine copulas build a hierarchical decomposition of the joint copula density by conditional factorisation into bivariate sub-components along dimensions (see Appendix B.2 and Aas et al., (2009); Czado, (2019); Czado and Nagler, (2022) for more detail). In practice, it is common to ignore part of the conditional relationships in an approach termed simplified vines (Nagler and Czado,, 2016; Nagler et al.,, 2017). This results in a convergence rate that is independent of the dimension so long as the true data-generating model is a simplified vine, and its correct decomposition has been identified (Nagler and Czado,, 2016). However, vines have an exponentially large model space with the dimension, meaning greedy search algorithms are employed in practice. Already for 5-dimensional data, itself simulated from a vine copula, recovering the correct decomposition with current vine models is extremely challenging (Janke et al.,, 2021), meaning these models are often misspecified in practice.
@@ 2.2 Probabilistic Classification for Density Ratio Estimation
The problem of density ratio estimation (DRE) is a crucial inference problem in machine learning and statistics (Sugiyama et al.,, 2010), most notably framed as a probabilistic classification task between samples from both densities involved in the ratio (Gutmann and Hyvärinen,, 2010; Gutmann and Hirayama,, 2011; Sugiyama et al.,, 2012; Menon and Ong,, 2016).
Given a dataset $\mathcal{D}_{p}=\{\mathbf{x}^{k}_{p}\}_{k=1}^{T}\sim p(\mathbf{x})$ and $\mathcal{D}_{q}=\{\mathbf{x}^{k}_{q}\}_{k=1}^{T}\sim q(\mathbf{x})$ with $q(\mathbf{x})>0$ whenever $p(\mathbf{x})>0$, we are interested in estimating the ratio $r(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$. Fictitious labels $y$ are assigned so that $y=1$ for $\mathcal{D}_{p}$ and $y=0$ for $\mathcal{D}_{q}$, with identical prior odds being assumed (although this can be relaxed). Then a probabilistic classifier $h(\mathbf{x}):\mathcal{X}\rightarrow[0,1]$ is trained to assign the probability of $\mathbf{x}$ being simulated from $p(\mathbf{x})$. By Bayes’ rule, the optimal value of $h(\mathbf{x})$ is $h^{*}(\mathbf{x})=\mathbb{P}(y=1|\mathbf{x})$, known as a Bayes optimal classifier (Hastie et al.,, 2009; Sugiyama et al.,, 2012; Gutmann and Hyvärinen,, 2012). We denote $\frac{h(\mathbf{x})}{1-h(\mathbf{x})}$ as the decision boundary since if this quantity is above one, it indicates $\mathbf{x}$ is more likely to belong to class $y=1$ than $y=0$. Crucially, the Bayes classifier can be equated to a density ratio via the decision boundary as follows:
@@
Developing methods for such problems is an active area of research, with new models being proposed for increasingly complex tasks (Rhodes et al.,, 2020; Choi et al.,, 2021, 2022).
@@ 3 Probabilistic Classification for Copula Estimation
In what follows, after establishing the connection between copula estimation and probabilistic classification, we derive a novel ratio copula estimator. Throughout, we assume that marginal densities are known and kept fixed, thus for $\mathbf{x}\sim p(\mathbf{x})$, we treat $\mathcal{P}(\mathbf{x})$ as samples from the true copula. This is a common approach in copula estimation (Fermanian,, 2005; Genève,, 2003; Scaillet,, 2007), and ensures the validity of the resulting copula-based density (Ashok et al., (2024) Proposition 2) even in non-parametric settings.
@@ 3.1 Ratio Copula Estimator
We propose to train a ratio copula estimator $c^{\text{ratio}}$ for the true copula density function $c$ using probabilistic classification. However, estimating the Bayes optimal classifier on $\mathcal{X}$ might not be straightforward depending on the constraints of the space, and estimating it on $[0,1]^{d}$ directly is even more cumbersome due to the constrained support. For this reason, we use the following transformation to $\mathbb{R}^{d}$, commonly used for copula models (Geenens et al.,, 2017)):
@@
Thus, we can now focus on the ratio of the RHS in (7) to train our classifier. Namely, our aim becomes training a classifier to distinguish datasets $\mathcal{D}_{\tilde{p}}=\Phi^{-1}\circ\mathcal{P}(\mathcal{D}_{p})$ from $\mathcal{D}_{\tilde{q}}=\Phi^{-1}\circ\mathcal{P}(\mathcal{D}_{\prod_{i}p_{i}})$, with respective labels $y=1$ and $y=0$, all conveniently on $\mathbb{R}^{d}$.
@@
where $\nu=\mathbb{P}(y=0)/\mathbb{P}(y=1)$ to compensate for unequal prior odds. The ratio copula $c^{\text{ratio}}(\cdot;\boldsymbol{\theta})$ implicitly defines a classifier between $\mathcal{D}_{\tilde{p}}$ and $\mathcal{D}_{\tilde{q}}$. As minimising $\mathcal{L}(\boldsymbol{\theta})$ is equivalent to minimising a Bregman divergence between $\mathcal{C}$ and ${c}(\cdot)$ (Menon and Ong,, 2016), $\exists\ \theta^{*}$ such that ${c^{\text{ratio}}}(\cdot;\boldsymbol{\theta}^{*})$ is the projection of $c(\cdot)$ on our model class $\mathcal{C}$. If our model class is well specified, then $c^{\text{ratio}}(\cdot;\theta^{*})=c(\cdot)$. In such a case, our estimator will asymptotically recover ${c^{\text{ratio}}}(\cdot;\boldsymbol{\theta}^{*})$, without needing any normalising constant (Gutmann and Hyvärinen,, 2012). However, as this only holds at optimality, we simultaneously estimate a normalising constant as an additional parameter $Z$ included in $\boldsymbol{\theta}$, which pre-multiplies the ratio copula. The estimation of $Z$ is done jointly with $\boldsymbol{\theta}$ by minimising the same loss, ensuring that the resulting ratio copula is a valid density.
Our ratio copula estimator inherits the theoretical guarantees associated with the loss in (A.1) from Gutmann and Hyvärinen, (2012), which we concretise here with a justification of its applicability given in Appendix A.1.
@@
Estimator properties; Gutmann and Hyvärinen, (2012) Assuming continuity of all marginal densities, ${c^{\text{ratio}}}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x});\boldsymbol{\theta}^{*})$ obtained by minimising Equation (A.1) is unique. Further, under certain regularity conditions detailed in Appendix A.1, the copula estimator is consistent with $\boldsymbol{\hat{\theta}}\overset{\mathbb{P}}{\to}\boldsymbol{\theta}^{*}$, and $\sqrt{|\mathcal{D}_{p}|}(\boldsymbol{\hat{\theta}}-\boldsymbol{\theta}^{*})$ is asymptotically normal.
Following the analysis from Gutmann and Hyvärinen, (2012) (Corollaries 4-7) of the loss function (A.1), our copula estimator’s accuracy is influenced by two main parts: the factor of independent to dependent observations $\nu$ and the form of $\tilde{q}$. Specifically, we want as many samples from $\tilde{q}$ as possible in order to inflate $\nu$, and we want $\tilde{q}$ to be similar to $\tilde{p}$. However, in our case, only dependent observations $\mathcal{D}_{\tilde{p}}$ are initially available and the form of $\tilde{q}$ is necessarily fixed by virtue of estimating a copula.

### v2

@@ 1 INTRODUCTION
Figure 1: Our ratio copula model estimates copula densities through classification. First panel: dependent copula data. Second panel: we transform the dependent copula data to a Gaussian scale and train a classifier to distinguish between dependent data and an independent Gaussian noise. Third panel: this classifier learns to identify dependence and ultimately recovers the copula density, shown on the log scale above.
Copulas are a popular tool for modelling multivariate densities by decomposing the estimation procedure into two steps: first, learning the marginal density of each variable; and second, stitching those marginals together through a copula model (the word copula comes from Latin and means “a link, tie, bond”). The adoption of copulas is primarily due to their distinct advantages over other models as a consequence of this two-step procedure. Namely, marginal models can be purpose-built for each dimension, which is relevant for applications, e.g. in neuroscience [Berkes et al., 2008], climate research [Huk et al., 2023], econometrics [Czado et al., 2022], and domain adaptation [Lopez-Paz et al., 2012]. Further, copulas focus all optimisation efforts uniquely on modelling dependence, leading to higher quality densities and samples [Tagasovska et al., 2019], and allows for the estimation of marginal models in parallel or in real time as used in Bayesian density estimation [Huk et al., 2024], Bayesian Networks [Elidan, 2010], and reinforcement learning [Wang et al., 2021].
@@
To this aim, we leverage a novel connection between copula densities and classification to propose a new copula estimation framework. This involves training a classifier to distinguish observed dependent data from an independent version of itself, recovering the true copula density at optimality. After providing a background on relevant topics in Section 2, we elicit the connection between copulas and classification in Section 3 and develop our classification-based copula estimation scheme. We show equivalences between copulas classifications in Section 4. We empirically validate our approach in Section 5, and discuss related works in Section 6 with our conclusion in Section 7.
@@ 2 BACKGROUND
Let $p(\mathbf{x})$ be a multivariate density over $\mathcal{X}\subseteq\mathbb{R}^{d}$, from which we observe i.i.d. samples $\mathcal{D}_{P}=\{\mathbf{x}^{k}\}_{k=1}^{T}\sim p(\mathbf{x})$. Similarly, let $p_{1}(x_{1}),\ldots,p_{d}(x_{d})$ be the marginal densities of $p(\mathbf{x})$, over $\mathcal{X}_{i}\subseteq\mathbb{R}$ with corresponding i.i.d. samples $\mathcal{D}_{P_{i}}=\{{x_{i}}^{k}\}_{k=1}^{T}\sim p_{i}({x_{i}})$ for $i=1,\ldots,d$, and assume that all distributions are continuous. Further, let $P$ and $P_{1},\ldots,P_{d}$ be the respective cumulative distribution functions (CDFs) of the previously mentioned densities, in order of appearance. We define applications of CDFs such as $\mathcal{P}(\mathbf{x})\equiv(P_{1}(x_{1}),\ldots,P_{d}(x_{d}))$ as shorthand for the dimension-wise application of marginal CDFs to a vector or $\mathcal{P}(\mathcal{D})$ to mean their application to every vector element of a set. Probability and cardinality of a set $A$ are denoted by $\mathbb{P}(A)$ and $|A|$ correspondingly.
@@ 2.1 Copula-Based Density Estimation
Sklar’s theorem (Sklar [1959] found in Appendix B) states that any joint probability density function with continuous marginal densities can always be uniquely factorised into the product of its marginals and a copula density $c:[0,1]^{d}\mapsto(0,\infty)$ in the following way:
@@
Consequently, one can decompose the estimation problem of learning $p(\mathbf{x})$ into first learning all marginals $\{p_{i}\}_{i=1}^{d}$, and then inferring an appropriate copula model for the dependence structure of $\mathcal{P}(\mathbf{x})\in[0,1]^{d}$.
For example, a scientist wishes to model daily temperature data over thousands of stations in different parts of a continent. With their domain expertise, they can design an effective model $p_{i}(x_{i}|S_{i})$ for the temperature $x_{i}$ of a given station $i$, taking into account relevant information $S_{i}$ such as the past days’ observed temperature and other variables such as rainfall, cloud cover and humidity in the area. Such a marginal model $p_{i}$ can be constructed for each station separately as it becomes a unidimensional time-series model with predictors. However, modelling all locations jointly while conditioning on every predictor across time and space might be unmanageable. In such cases, a copula is the appropriate tool to stitch the separate station-specific models together into a mathematically valid joint model over the whole space using (2.1).
Whereas taking the product of marginal models $\prod p_{i}(x_{i})$ would also give a valid density model, the joint behaviour would be unaccounted for (represented as an independence copula giving value 1 to all possible observations $\mathcal{P}(\mathbf{x})$), effectively assuming independence across all stations. This is an issue since, for instance, the temperature at two nearby stations should be similar - more than that of two stations located further apart. This is exactly the joint behaviour that a copula serves to capture.
@@
There exist many types of copula models, both parametric and non-parametric. Most parametric copulas are only really suited for modelling 2D dependences and greatly suffer from the curse of dimensionality [Hofert et al., 2013]. The Gaussian copula (see Appendix B.1) is a popular parametric choice as it can be fitted quickly even to moderate dimensions. However, it lacks the desired flexibility to capture complex patterns or heavy tails; one of the argued reasons for the 2008 financial crash [Salmon, 2012]. Archimedian and Archimax parametric copula families [Ng et al., 2022, Górecki and Okhrin, 2024] have similar drawbacks. Among non-parametric copulas, kernel density estimator (KDE) copulas [Geenens et al., 2017] are commonly used. They transform the observed data via appropriate (often Gaussian) inverse CDFs to a latent space and perform regular KDE on the latent density. However, doing so suffers from drawbacks common to KDE methods in higher dimensions. Deep learning copula models are still a nascent line of research, with a handful of candidate solutions such as Kamthe et al. [2021], Janke et al. [2021], Ashok et al. [2024]. The Implicit Generative Copula (IGC) of Janke et al. [2021] is a promising model, which works by generating data in a latent space and applying an approximate integral transformation to recover samples in $[0,1]$. As an implicit model only able to generate samples, the IGC cannot evaluate densities, thus preventing many applications of copula models. As such, current copula models supporting density estimation and sampling are mostly limited to low or medium-dimensional modelling.
@@
The state-of-the-art approach is vine copula models, offering a divide-and-conquer solution to expand copula models to higher dimensions. Vine copulas build a hierarchical decomposition of the joint copula density by conditional factorisation into bivariate sub-components along dimensions (see Appendix B.2 and Aas et al. [2009], Czado [2019], Czado and Nagler [2022] for more detail). In practice, it is common to ignore part of the conditional relationships in an approach termed simplified vines [Nagler and Czado, 2016, Nagler et al., 2017]. This results in a convergence rate that is independent of the dimension so long as the true data-generating model is a simplified vine, and its correct decomposition has been identified [Nagler and Czado, 2016]. However, vines have an exponentially large model space with the dimension, meaning greedy search algorithms are employed in practice. Already for 5-dimensional data, itself simulated from a vine copula, recovering the correct decomposition with current vine models is extremely challenging [Janke et al., 2021], meaning these models are often misspecified in practice.
@@ 2.2 Probabilistic Classification for Density Ratio Estimation
The problem of density ratio estimation (DRE) is a crucial inference problem in machine learning and statistics [Sugiyama et al., ], most notably framed as a probabilistic classification task between samples from both densities involved in the ratio [Gutmann and Hyvärinen, 2010, Gutmann and Hirayama, 2011, Sugiyama et al., 2012, Menon and Ong, 2016].
Given a dataset $\mathcal{D}_{p}=\{\mathbf{x}^{k}_{p}\}_{k=1}^{T_{p}}\sim p(\mathbf{x})$ and $\mathcal{D}_{q}=\{\mathbf{x}^{k}_{q}\}_{k=1}^{T_{q}}\sim q(\mathbf{x})$ with $q(\mathbf{x})>0$ whenever $p(\mathbf{x})>0$, we are interested in estimating the ratio $r(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$. Fictitious labels $y$ are assigned so that $y=1$ for $\mathcal{D}_{p}$ and $y=0$ for $\mathcal{D}_{q}$, with identical prior odds being assumed (although this can be relaxed). Then a probabilistic classifier $h(\mathbf{x}):\mathcal{X}\rightarrow[0,1]$ is trained to assign the probability of $\mathbf{x}$ being simulated from $p(\mathbf{x})$. By Bayes’ rule, the optimal value of $h(\mathbf{x})$ is $h^{*}(\mathbf{x})=\mathbb{P}(y=1|\mathbf{x})$, known as a Bayes optimal classifier [Hastie et al., 2009, Sugiyama et al., 2012, Gutmann and Hyvärinen, 2012]. We denote $\frac{h(\mathbf{x})}{1-h(\mathbf{x})}$ as the decision boundary since if this quantity is above one, it indicates $\mathbf{x}$ is more likely to belong to class $y=1$ than $y=0$. Crucially, the Bayes classifier can be equated to a density ratio via the decision boundary as follows:
@@
Developing methods for such problems is an active area of research, with new models being proposed for increasingly complex tasks [Rhodes et al., 2020, Choi et al., 2021, 2022].
@@ 3 PROBABILISTIC CLASSIFICATION FOR COPULA ESTIMATION
In what follows, after establishing the connection between copula estimation and probabilistic classification, we derive a novel ratio copula estimator. Throughout, we assume that marginal densities are known and kept fixed, thus for $\mathbf{x}\sim p(\mathbf{x})$, we treat $\mathcal{P}(\mathbf{x})$ as samples from the true copula. This is a common approach in copula estimation [Fermanian, 2005, Genève, 2003, Scaillet, 2007], and ensures the validity of the resulting copula-based density (Ashok et al. [2024] Proposition 2) even in non-parametric settings.
@@ 3.1 Ratio Copula Estimator
We propose to train a ratio copula estimator $c^{\text{ratio}}$ for the true copula density function $c$ using probabilistic classification. However, estimating the Bayes optimal classifier on $\mathcal{X}$ might not be straightforward depending on the constraints of the space, and estimating it on $[0,1]^{d}$ directly is even more cumbersome due to the constrained support. For this reason, we use the probit transformation $\Phi^{-1}:[0,1]\mapsto\mathbb{R}$ applied dimension-wise, commonly used for copula models [Geenens et al., 2017]):
@@
Thus, we can now focus on the ratio of the RHS in (7). Namely, our aim becomes training a classifier to distinguish dependent data $\mathcal{D}_{\tilde{p}}=\Phi^{-1}\circ\mathcal{P}(\mathcal{D}_{p})$ from independent data $\mathcal{D}_{\tilde{q}}=\Phi^{-1}\circ\mathcal{P}(\mathcal{D}_{\prod_{i}p_{i}})$, with respective labels $y=1$ and $y=0$, all conveniently on $\mathbb{R}^{d}$.
@@
where $\nu=\mathbb{P}(y=0)/\mathbb{P}(y=1)$ to compensate for unequal prior odds. The ratio copula $c^{\text{ratio}}(\cdot;\boldsymbol{\theta})$ implicitly defines a classifier between $\mathcal{D}_{\tilde{p}}$ and $\mathcal{D}_{\tilde{q}}$. As minimising $\mathcal{L}(\boldsymbol{\theta})$ is equivalent to minimising a Bregman divergence between $\mathcal{C}$ and ${c}(\cdot)$ [Menon and Ong, 2016], $\exists\ \theta^{*}$ such that ${c^{\text{ratio}}}(\cdot;\boldsymbol{\theta}^{*})$ is the projection of $c(\cdot)$ on our model class $\mathcal{C}$. If our model class is well specified, then $c^{\text{ratio}}(\cdot;\theta^{*})=c(\cdot)$. In such a case, our estimator will asymptotically recover ${c^{\text{ratio}}}(\cdot;\boldsymbol{\theta}^{*})$, without needing any normalising constant [Gutmann and Hyvärinen, 2012]. However, as this only holds at optimality, we simultaneously estimate a normalising constant as an additional parameter $Z$ included in $\boldsymbol{\theta}$, which pre-multiplies the ratio copula. The estimation of $Z$ is done jointly with $\boldsymbol{\theta}$ by minimising the same loss, ensuring that the resulting ratio copula is a valid density.
Our ratio copula estimator inherits the theoretical guarantees associated with the loss in (3.1) from Gutmann and Hyvärinen [2012], which we concretise here with a justification of its applicability given in Appendix A.1.
@@
Estimator properties; Gutmann and Hyvärinen [2012] Assuming continuity of all marginal densities, ${c^{\text{ratio}}}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x});\boldsymbol{\theta}^{*})$ obtained by minimising Equation (3.1) is unique. Further, under certain regularity conditions detailed in Appendix A.1, the copula estimator is consistent with $\boldsymbol{\hat{\theta}}\overset{\mathbb{P}}{\to}\boldsymbol{\theta}^{*}$, and $\sqrt{|\mathcal{D}_{p}|}(\boldsymbol{\hat{\theta}}-\boldsymbol{\theta}^{*})$ is asymptotically normal.
Following the analysis from Gutmann and Hyvärinen [2012] (Corollaries 4-7) of the loss function (3.1), our copula estimator’s accuracy is influenced by two main parts: the factor of independent to dependent observations $\nu$ and the form of $\tilde{q}$. Specifically, we want as many samples from $\tilde{q}$ as possible in order to inflate $\nu$, and we want $\tilde{q}$ to be similar to $\tilde{p}$. However, in our case, only dependent observations $\mathcal{D}_{\tilde{p}}$ are initially available and the form of $\tilde{q}$ is necessarily fixed by virtue of estimating a copula.